# MultiConIR: Towards multi-condition Information Retrieval

### v2

@@ Abstract
In this paper, we introduce MultiConIR, the first benchmark designed to evaluate retrieval models in multi-condition scenarios. Unlike existing datasets that primarily focus on single-condition queries from search engines, MultiConIR captures real-world complexity by incorporating five diverse domains: books, movies, people, medical cases, and legal documents. We propose three tasks to systematically assess retrieval and reranking models on multi-condition robustness, monotonic relevance ranking, and query format sensitivity. Our findings reveal that existing retrieval and reranking models struggle with multi-condition retrieval, with rerankers suffering severe performance degradation as query complexity increases. We further investigate the performance gap between retrieval and reranking models, exploring potential reasons for these discrepancies, and analysis the impact of different pooling strategies on condition placement sensitivity. Finally, we highlight the strengths of GritLM and Nv-Embed, which demonstrate enhanced adaptability to multi-condition queries, offering insights for future retrieval models. The code and datasets are available at https://github.com/EIT-NLP/MultiConIR
@@
Xuan Lu1,2, Sifan Liu2, Bochao Yin2, Yongqi Li3, Xinghao Chen2,3, Hui Su4, Yaohui Jin1, Wenjun Zeng2, Xiaoyu Shen2thanks: Corresponding author, 1Shanghai Jiao Tong University 2Ningbo Key Laboratory of Spatial Intelligence and Digital Derivative, Institute of Digital Twin, EIT 3Department of Computing, The Hong Kong Polytechnic University 4Meituan Inc. xyshen@eitech.edu.cn
@@ 1 Introduction
Information retrieval (IR) is critical for helping users find relevant information across various domains. Traditionally, IR systems retrieve documents by matching queries based on lexical similarity, such as BM25 (Carpineto and Romano, 2012; Ponte and Croft, 2017), or semantic similarity using dense vector representations (Karpukhin et al., 2020; Zhan et al., 2021). Though highly effective for queries with straightforward query-document relationships (Su et al., 2024), they often fail to fully capture nuanced intent as user needs become more complex (Zhu et al., 2023; Su et al., 2024).
Figure 1: From single-condition to multi-condition retrieval. Standard retrieval relies on single-condition queries, while instruction-following retrieval enhances queries with explicit instructions. Traditional SQL-based filtering handles predefined attributes, whereas real-world multi-condition retrieval involves queries specifying multiple conditions, which can be attribute-level or semantic-level.
@@
A significant challenge arises when users specify multiple requirements simultaneously, as illustrated in Fig 1. Whether searching for a movie with specific attributes or selecting a product that meets various criteria, multi-condition search has become an integral part of modern information-seeking behavior. Traditional IR systems handle such scenarios using structured filtering, such as SQL-based queries that retrieve information from backend databases based on predefined conditions. However, this approach is inherently rigid and limited, as it relies on explicitly defined attributes and lacks the flexibility to accommodate evolving or diverse user preferences. As a result, it struggles to support nuanced or semantic-level queries that go beyond structured data filtering.
@@
The advent of Large Language Models (LLMs) has enhanced IR by introducing instruction-following capabilities (Asai et al., 2023; Weller et al., 2024a; Oh et al., 2024). This approach augments standard queries with explicit instructions, which serve as additional constraints to refine search results, as shown in Fig.1. Despite these advancements, existing evaluation benchmarks remain predominantly focused on single-condition queries and binary relevance assessments—classifying documents as either relevant or irrelevant (Nguyen et al., 2016; Kwiatkowski et al., 2019; Muennighoff et al., 2022)—thus overlooking the nuanced challenges of multi-condition queries, where relevance depends on the degree to which multiple conditions are satisfied.
@@
An ideal multi-condition retrieval system should exhibit the following properties: (1) Complexity Robustness: The system should maintain high performance regardless of query complexity (i.e., the number of conditions specified); (2) Relevance Monotonicity: The relevance scores should scale monotonically with the number of matched conditions; for example, a document matching all $n$ conditions should be ranked higher than one matching $n-1$; (3) Format Invariance: The system should yield consistent results regardless of the query format, whether presented as a structured list or as free-form natural language.
@@
Existing benchmarks do not offer a structured framework for evaluating multi-condition retrieval along these dimensions. To address this gap, we introduce MultiCondIR—the first benchmark designed to comprehensively evaluate multi-condition retrieval systems. Through systematic experiments on 12 state-of-the-art models (including dense retrievers, cross-encoders, and LLM-based agents), we uncover several critical insights:
  • Multi-Condition Degradation: Dense retrieval and reranking models deteriorate as query conditions increase, with rerankers suffering more severely. Rerankers’ token-level interactions benefit single-condition tasks but hinder multi-condition retrieval.
  • Relevance Monotonicity Failure: Models fail to maintain a consistent ranking hierarchy. Dense retrievers tend to focus more on “exact match” and “complete mismatch”, likely due to their training paradigm.
  • Query Format Sensitivity: Models are susceptible to query phrasing, with rerankers exhibiting a flip rate exceeding 20%.
  • Pooling Impact on Focus: Different embedding pooling strategies influence the model’s focus on query condition positions. Mean pooling focuses on early query conditions, <EOS> pooling biases toward later conditions, while latent attention layer pooling distributes focus more evenly.
  • GritLM’s Advantage: GritLM outperforms other models, potentially due to its hybrid attention mechanism, which integrates both bidirectional and causal attention to improve the understanding of complex queries.
@@
By quantifying these gaps, our work provides actionable guidelines for developing IR systems that truly understand multi-condition intent, laying the groundwork for advancing IR toward human-like reasoning in complex search scenarios.
@@ 2 Related Works
Traditional sparse retrieval methods are based on BM25 (Robertson and Zaragoza, 2009), TF-IDF (Ramos et al., 2003), etc., rely on keyword matching and statistical weighting to evaluate relevance, which suffers from the well-known issue of lexical gap (Berger et al., 2000), restricting their ability to effectively capture semantic relationships (Luan et al., 2021; Nian et al., 2024).
Dense retrieval addresses this limitation by encoding both queries and documents as embeddings within a joint latent space, where the semantic relationship is captured through the similarity scores between their embeddings (Li et al., 2023a). Pre-trained language models like BERT (Devlin et al., 2019) and T5 (Raffel et al., 2020), are widely used as backbone encoders for dense retrieval (Li et al., 2023b; Sturua et al., 2024; Xiao et al., 2023).
Recent advancements have shown that LLMs offer significant potential as backbone encoders for dense retrieval (Wang et al., 2024a; Weller et al., 2024c; BehnamGhader et al., 2024). For instance, Repllama (Ma et al., 2023) enhanced retrieval performance by fine-tuning Llama-2 to serve as dense retrievers. GritLM (Muennighoff et al., 2024) unified text embedding and generation within a single LLM. LLM2Vec (BehnamGhader et al., 2024) introduced an unsupervised approach for transforming decoder-only LLMs into dense retrievers.
@@
Existing datasets for information retrieval, such as MS MARCO (Nguyen et al., 2016), Natural Questions (Kwiatkowski et al., 2019), and MTEB (Muennighoff et al., 2022), primarily focus on queries sourced from search engines. The relationships between queries and documents are typically simple and direct (Su et al., 2024).
Recent studies have expanded retrieval benchmarks to address more complex scenarios. Instruction-based datasets (Weller et al., 2024a; Qin et al., 2024; Oh et al., 2024), for instance, evaluate the instruction-following capabilities of retrieval models by embedding explicit instructions within queries to better represent users’ retrieval intents. Furthermore, some works have assessed retrieval models’ abilities to handle logical reasoning tasks, including Boolean logic Mai et al. (2024); Zhang et al. (2024c), negation Zhang et al. (2024a); Weller et al. (2024b), and multi-hop reasoning Su et al. (2024). These efforts mark significant progress in increasing query complexity. However, while research in the generative modeling domain has explored the ability of LLMs to handle multi-constraint instructions (He et al., 2024; Ferraz et al., 2024; Zhang et al., 2024b), studies on retrieval models in multi-condition scenarios remain sparse.
@@ 3 MultiConIR
(3) A sequence of hard negative documents $\{d_{0},d_{1},\dots,d_{k-1}\}$, where each $d_{j}$ satisfies exactly $j$ out of $k$ conditions, formally expressed as $d_{j}\models\{c_{1},\dots,c_{j}\}$ and $d_{j}\not\models\{c_{j+1},\dots,c_{k}\}$.
@@ 3.1 Domain Selection
MultiConIR was constructed from real-world datasets from five diverse domains: books, movies, people, medical cases, and legal documents. These domains were carefully chosen for their practical significance and their intrinsic requirement for multi-condition retrieval. Books & Movies: Queries in these domains often blend attribute-based (e.g., a sci-fi movie) with narrative-based (e.g., a story about time travel), requiring semantic understanding beyond keyword matching. People: Queries are often based on vague recollections or specific but incomplete information about a person (e.g., a Nobel laureate in Physics who studied black holes) Medical Case & Legal Document: These domains require fine-grained condition matching, where case-based reasoning is essential (e.g., a patient with breathing issues and antibiotic allergies), demanding models understand condition dependencies, not just match isolated terms.
@@ 3.2 Dataset Construction Pipeline
Figure 2: MultiCon Dataset Construction Pipeline.
@@
To construct MultiConIR, we design a multi-step data generation framework as shown in Fig. 2, this pipeline is highly adaptable across multiple domains, enabling the generation of queries and hard negative (HN) documents that progressively satisfy 1 to 10 conditions. To preserve dataset integrity and mitigate the generalization issues associated with fully synthetic datasets (Li et al., 2023c; Wang et al., 2024b), we employ LLM-based generation (GPT-4o) exclusively for modifying sentences within hard negatives, rather than altering entire documents. [^1] Discussion about the challenges associated with fully LLM-generated datasets can be found in Appendix E.1. The data creation process consists of the following steps, with detailed prompt templates provided in Appendix B:
Step 1: Condition Sentence Extraction. To capture the fine-grained constraints for multi-condition retrieval, we prompt GPT-4o to extract ten key sentences from each document, where each sentence represents a distinct and semantically complete condition. To ensure robustness, extracted sentences must be contextually relevant, non-redundant, and representative of the document’s key information.
@@
Step 2: Query Generation. Given the extracted condition sentences, we prompt GPT-4o to generate ten hierarchical queries (Query1 to Query10), each incorporating an increasing number of conditions (from 1 to 10). To enhance linguistic diversity, queries are formulated in two distinct styles: (1) Instruction-style (explicitly listing conditions in a structured format). (2) Descriptive-style (embedding conditions naturally within a coherent sentence).
Step 3: Hard Negative Sentence Construction. To create hard negatives, we prompt GPT-4o to modify extracted condition sentences into Hard Negative Sentences (HNS) that subtly deviate from the original meaning while preserving lexical plausibility. Unlike naive negation, these modifications introduce semantic shifts that retain fluency but fail to satisfy the original retrieval conditions, thereby making retrieval harder. We experimented with two approaches for HNS construction: one modifies key information (applied to books, movies, medical cases, and legal documents), while the other retains keywords but adds dummy information (used for the people dataset).  [^2] The differences between these two approaches are discussed in Appendix E.2.
Step 4: HN Document Generation. To construct multi-condition retrieval datasets, we systematically assemble documents: (1) Positive documents: Composed of all ten original condition sentences, ensuring full query relevance. (2) Hard Negative documents $d_{1}$ to $d_{10}$: Generated by progressively replacing 1 to 10 original condition sentences with their corresponding hard negative versions, creating a continuum of relevance degradation. This structured data generation ensures that each query condition incrementally refines document ranking, allowing for fine-grained retrieval performance assessment.

### v3

@@ Abstract
Multi-condition information retrieval (IR) presents a significant, yet underexplored challenge for existing systems. This paper introduces MultiConIR, a benchmark specifically designed to evaluate retrieval and reranking models under nuanced multi-condition query scenarios across five diverse domains. We systematically assess model capabilities through three critical tasks: complexity robustness, relevance monotonicity, and query format sensitivity. Our extensive experiments on 15 models reveal a critical vulnerability: most retrievers and rerankers exhibit severe performance degradation as query complexity increases. Key deficiencies include widespread failure to maintain relevance monotonicity, and high sensitivity to query style and condition placement. The superior performance of GPT-4o reveals the performance gap between IR systems and advanced LLM for handling sophisticated natural language queries. Furthermore, this work delves into the factors contributing to reranker performance deterioration and examines how condition positioning within queries affects similarity assessment, providing crucial insights for advancing IR systems towards complex search scenarios. The code and datasets are available at https://github.com/EIT-NLP/MultiConIR
@@
Xuan Lu1,2,  Sifan Liu2thanks: Equal contribution,  Bochao Yin2footnotemark: ,  Yongqi Li3,  Xinghao Chen2,3, Hui Su4,  Yaohui Jin1,  Wenjun Zeng2,  Xiaoyu Shen2thanks: Corresponding author 1Shanghai Jiao Tong University 2Ningbo Key Laboratory of Spatial Intelligence and Digital Derivative, Institute of Digital Twin, EIT 3Department of Computing, The Hong Kong Polytechnic University 4Meituan Inc. lux1997@sjtu.edu.cn xyshen@eitech.edu.cn
@@ 1 Introduction
Information retrieval (IR) is critical for helping users find relevant information across various domains. Traditionally, IR systems retrieve documents by matching queries based on lexical similarity, such as BM25 (Carpineto and Romano, 2012; Ponte and Croft, 2017), or semantic similarity using dense vector representations (Karpukhin et al., 2020; Zhan et al., 2021). Though highly effective for queries with straightforward query-document relationships (Su et al., 2024), they often fail to fully capture nuanced intent as user needs become more complex (Zhu et al., 2023; Su et al., 2024; Zhu et al., 2025).
Figure 1: From single-condition to multi-condition retrieval. Standard and instruction-aware retrieval address single-condition queries. SQL-based filtering is restricted to predefined attributes within structured data. Real-world multi-condition retrieval enables the formulation of multiple, often semantic, conditions using natural language query.
@@
A significant challenge arises when users specify multiple requirements simultaneously, as illustrated in Fig. 1. Whether searching for a movie with specific attributes or selecting a product that meets various criteria, multi-condition search has become an integral part of modern information-seeking behavior. Traditional IR systems handle such scenarios using structured filtering, such as SQL-based queries that retrieve information from backend databases based on predefined conditions. However, this approach is inherently rigid and limited, as it relies on explicitly defined attributes and lacks the flexibility to accommodate evolving or diverse user preferences. As a result, it struggles to support nuanced or semantic-level queries that go beyond structured data filtering.
@@
The advent of Large Language Models (LLMs) has enhanced IR by introducing instruction-following capabilities (Asai et al., 2023; Weller et al., 2024a; Oh et al., 2024; Liu et al., 2024). This approach augments standard queries with explicit instructions, which serve as additional constraints to refine search results, as shown in Fig.1. Despite these advancements, existing evaluation benchmarks remain predominantly focused on single-condition queries and binary relevance assessments—classifying documents as either relevant or irrelevant (Nguyen et al., 2016; Kwiatkowski et al., 2019; Muennighoff et al., 2022)—thus overlooking the nuanced challenges of multi-condition queries, where relevance depends on the degree to which multiple conditions are satisfied.
@@
An ideal multi-condition retrieval system should exhibit the following properties: (1) Complexity Robustness: The system should maintain high performance regardless of query complexity (i.e., the number of conditions specified); (2) Relevance Monotonicity: The relevance scores should scale monotonically with the number of matched conditions; for example, a document matching all $n$ conditions should be ranked higher than one matching n1n-1; (3) Format Invariance: The system should yield consistent results regardless of the query format, whether presented as a structured list or as free-form natural language.
@@
Existing benchmarks do not offer a structured framework for evaluating multi-condition retrieval along these dimensions. To address this gap, we introduce MultiConIR—a benchmark designed to comprehensively evaluate multi-condition retrieval systems. Through systematic experiments on 15 state-of-the-art models (including dense retrievers, cross-encoders, and LLM-based agents), we uncover several critical insights:
  • Multi-Condition Struggle: Retrievers and Rerankers struggle with multi-condition retrieval, showing performance decline as query conditions increase, difficulty with relevance monotonicity, and sensitivity to query style variations.
  • Retrievers and Rerankers Differ: Rerankers excel with single-condition queries but fail under multiple conditions. Retrievers demonstrate greater robustness. GritLM demonstrates the best robustness among retrievers.
  • Position Impacts Model Focus: Dense retriever pooling strategies emphasize different condition positions, mean pooling focuses on initial positions, while <EOS> pooling emphasizes final positions. Rerankers exhibit non-uniform attention across positions and greater sensitivity to document length variations.
@@
By quantifying these gaps, our work reveals key deficiencies in the ability of current IR systems to understand multi-condition intent, laying the groundwork for advancing IR toward human-like reasoning in complex search scenarios.
@@ 2 Related Works
Traditional sparse retrieval methods are based on BM25 (Robertson and Zaragoza, 2009), TF-IDF (Ramos et al., 2003), etc., rely on keyword matching and statistical weighting to evaluate relevance, which suffers from the well-known issue of lexical gap (Berger et al., 2000), restricting their ability to effectively capture semantic relationships (Luan et al., 2021; Nian et al., 2024). Dense retrieval addresses this limitation by encoding both queries and documents as embeddings within a joint latent space, where the semantic relationship is captured through the similarity scores between their embeddings (Li et al., 2023a). Pre-trained language models like BERT (Devlin et al., 2019) and T5 (Raffel et al., 2020), are widely used as backbone encoders for dense retrieval (Li et al., 2023b; Sturua et al., 2024; Xiao et al., 2023). Recent advancements have shown that LLMs offer significant potential as backbone encoders for dense retrieval (Wang et al., 2024a; Weller et al., 2024c; BehnamGhader et al., 2024). For instance, Repllama (Ma et al., 2023) fine-tuned Llama-2 to serve as dense retrievers. GritLM (Muennighoff et al., 2024) unified text embedding and generation within a single LLM. LLM2Vec (BehnamGhader et al., 2024) introduced an unsupervised approach for transforming decoder-only LLMs into dense retrievers.
@@
Existing datasets for information retrieval, such as MS MARCO (Nguyen et al., 2016), Natural Questions (Kwiatkowski et al., 2019), and MTEB (Muennighoff et al., 2022), primarily focus on queries sourced from search engines. The relationships between queries and documents are typically simple and direct (Su et al., 2024). Recent studies have expanded retrieval benchmarks to address more complex scenarios. BIRCO (Wang et al., 2024b) introduces a benchmark designed to evaluate information retrieval tasks with complex objectives. Instruction-based datasets (Weller et al., 2024a; Qin et al., 2024; Oh et al., 2024), for instance, evaluate the instruction-following capabilities of retrieval models by embedding explicit instructions within queries to better represent users’ retrieval intents. Furthermore, some works have assessed retrieval models’ abilities to handle logical reasoning tasks, including Boolean logic Mai et al. (2024); Zhang et al. (2024c), negation Zhang et al. (2024a); Weller et al. (2024b), and multi-hop reasoning Su et al. (2024); Liu et al. (2025). These efforts mark significant progress in increasing query complexity. However, while research in the generative modeling domain has explored the ability of LLMs to handle multi-constraint instructions (He et al., 2024; Ferraz et al., 2024; Zhang et al., 2024b), studies on retrieval models in multi-condition scenarios remain sparse.
@@ 3 MultiConIR
(3) A sequence of hard negative (HN) documents $\{d_{0},d_{1},\dots,d_{k-1}\}$, where each $d_{j}$ satisfies exactly $j$ out of $k$ conditions, formally expressed as $d_{j}\models\{c_{1},\dots,c_{j}\}$ and $d_{j}\not\models\{c_{j+1},\dots,c_{k}\}$.
@@ 3.1 Domain Selection
To construct the MultiConIR dataset, we meticulously selected five domains—Books, Movies, People, Medical Cases, and Legal Documents—each chosen for its practical significance and inherent suitability for evaluating multi-condition retrieval capabilities.
Books & Movies: These domains represent common consumer searches where nuanced preferences are expressed by combining structured attributes (e.g., genre, creator, year, cast) with narrative elements (e.g., plot details, thematic content like “a story about time travel”). Effective retrieval demands semantic understanding beyond simple keyword matching to process multifaceted queries, such as "an action film directed by Christopher Nolan, starring Leonardo DiCaprio, released after 2010, with an intense chase scene."
People: Queries about individuals frequently rely on partial or vague information, such as notable achievements or specific traits. An example query could be “a Nobel laureate in Physics who studied black holes.” These searches demand that IR systems effectively handle incomplete data and infer connections between various attributes to identify the correct individual.
Medical Case & Legal Document: The medical case and legal document domains offer more practical and application-driven use cases. In the medical domain, doctors often rely on retrieval systems to reference historical cases to support diagnostic decisions. A typical query might include multiple conditions, such as “Find a case that meets the following conditions: 1) middle-aged female patient; 2) hospitalized for breathing difficulties; 3) has a history of antibiotic allergies; 4) has a family history of peanut allergies.” Similarly, in the legal domain, retrieval users often seek case law with high similarity to ongoing cases, which requires matching various legal and factual attributes in historical court decisions. These complex queries require IR systems to perform fine-grained condition matching and understand the interdependencies between various factors.
@@ 3.2 Dataset Construction Pipeline
Figure 2: MultiConIR Dataset Construction Pipeline: (1) relevant condition sentences are extracted from documents; (2) these conditions are then used to generate multi-condition queries; (3) hard negative (HN) versions of the condition sentences are created; and (4) positive documents and progressively challenging HN documents are assembled from these elements.
@@
To construct MultiConIR, we design a multi-step data generation framework. As shown in Fig. 2, this pipeline is highly adaptable across multiple domains, enabling the generation of queries and hard negative (HN) documents that progressively satisfy 1 to 10 conditions. To preserve dataset integrity and mitigate the generalization issues associated with fully synthetic datasets (Li et al., 2023c; Wang et al., 2024c), we employ LLM-based generation (GPT-4o) exclusively for modifying sentences within hard negatives, rather than altering entire documents. [^1] Fully LLM-generated datasets may introducing inherent linguistic biases of the underlying LLMs, and lacks the contextual richness and complexity in real-world retrieval (Shumailov et al., 2024). To mitigate these issues, we restricted LLM interventions to modifying only condition sentences rather than entire document. We further discuss this problem in Appendix G.1. The data creation process consists of the following steps, with detailed prompt templates provided in Appendix B:
Step 1: Condition Sentences Extraction. For each source document $d$, we issue a structured prompt to GPT-4o that (1) reads the entire text, (2) identifies ten non-overlapping conditions that describe key features expressed in $d$, and (3) return ten original sentences, where every sentence is semantically complete, and expresses a distinct condition. Documents yielding fewer than ten qualified sentences are discarded. The final ten sentences constitute the condition set $C(d)$, which we later recombine into queries of varying complexity and use as the ground-truth positive for multi-condition retrieval experiments.
@@
Step 2: Query Generation. Given the ten-sentence condition set $C(d)$, we prompt GPT-4o to synthesise a hierarchy of ten queries $\{q_{1},q_{2},...,q_{10}\}$. Each $q_{k}$ incorporating the first $k$ conditions. To enhance linguistic diversity, For every $q_{k}$ we request two forms: (1) Instruction-style $q_{k}^{\text{inst}}$: a bullet-like template (“Find a document that satisfies: (1)… (2)…”) for explicit parsing. (2) Descriptive-style $q_{k}^{\text{des}}$: embedding conditions naturally within a coherent sentence.
Step 3: Hard Negative Sentence Construction. For each condition sentence $c_{i}\in C(d)$ we instruct GPT-4o to produce a semantically divergent yet fluently written variant $h_{i}$ that no longer satisfies the original constraint. The rewrite must preserve overall length and style while introducing either (i) a subtle alteration of a critical fact (applied to books, movies, medical cases, and legal documents) or (ii) an innocuous clause that injects misleading information without changing the existing keywords (used for the people corpus).  [^2] Retrievers are more robust when adding misleading information; Modifying critical facts is challenging for both retrievers and rerankers. A detailed comparison of these two strategies is given in Appendix G.2. The ten variants form the hard-negative set $\mathrm{HNS}(d)=\{h_{1},\dots,h_{10}\}$.
Step 4: Hard Negative Document Generation. Starting from the ten-sentence condition set $C(d)=\{c_{1},\dots,c_{10}\}$ and its hard-negative counterparts $\mathrm{HNS}(d)=\{h_{1},\dots,h_{10}\}$, we build an ordinal ladder of document variants. The positive document $d^{+}$ contains the full sequence $c_{1},\dots,c_{10}]$. For each $k\in\{1,\dots,10\}$ we generate a HN document $d_{k}^{-}=\bigl{[}h_{1},\dots,h_{k},\;c_{k+1},\dots,c_{10}\bigr{]},$ i.e., the first $k$ conditions are replaced by their semantically perturbed versions while the remaining 10k10-k conditions stay intact. This yields a controlled degradation chain $d^{+}\!\rightarrow d_{1}^{-}\!\rightarrow\cdots\!\rightarrow d_{10}^{-}$, ranging from a single violated constraint to a completely adversarial variant. Coupled with the hierarchical queries $\{Q_{k}\}_{k=1}^{10}$, the corpus enables fine-grained evaluation of retrieval performance under progressively stricter condition sets.
Figure 3: Benchmark quality evaluation framework of MultiConIR. Query realism was assessed by human annotators. Label accuracy involved initial GPT-4o filtering, followed by a final human double-check.
Benchmark Quality Assurance.
The reliability of MultiConIR was audited on two complementary fronts (Fig. 3). (1) Query validity. From each of the five domains we randomly sampled 100 multi-condition queries (500 in total) and asked ten trained annotators to judge whether each query was natural, precise, and contextual plausibility (valid vs. unvalid). The resulting inter-annotator agreement reached $93.7\%$, with Fleiss’ $\kappa=0.84$. (2) Document–label validity. To detect false positives/negatives, we first applied GPT-4o to the entire corpus: the model verified for every document variant $d_{k}^{-}$ whether exactly $k$ of the ten conditions were satisfied, and we discarded mismatched instances. We then drew another 100 document–query pairs per domain (500 total) for manual spot-checking; two independent annotators reviewed each pair and a third adjudicated disagreements, yielding a residual error rate of $2.4\%$. After these filtering steps the final benchmark sizes for each domain are summarised in Table 1, and the full evaluation protocol is detailed in Appendix C.
Domain Number Avg. Length Source License Example
$D^{+}$ $Q$ $D^{-}$ $Q$ $D$
People 420 4200 4200 31.2 225.2 People Wikipedia Dat(Mahajan, 2017) CC0 1.0 Table 8
Books 482 4820 4820 25.6 235.3 Books Dataset (Rustamov, 2021) CC0 1.0 Table 9
Movies 500 5000 5000 24.8 184.6 Wikipedia Movie Plots (Robischon, 2018) CC BY-SA 4.0 Table 10
Medical Case 479 4790 4790 28.4 212.1 Medical Transcription (MTSamples, 2025) CC0 1.0 Table 11
Legal Document 426 4260 4260 34.1 302.2 LexGLUE (Chalkidis et al., 2022) CC BY 4.0 Table 12
Table 1: Data statistics of MutiConIR. For each dataset, we show the number of positive documents ($D^{+}$ ), queries ($Q$) and hard negative documents($D^{-}$), and the average length (in words) of queries and documents, and the source dataset of each domain.